During a 6sec time interval a fly wheel with a constant angular acceleration turns through 500 radians that acquire an angular velocity of 100rad/sec.
A, what is the angular velocity at the beginning of the 6 sec?
B, what is the acceleration of the fly wheel?
1
Expert's answer
2020-05-07T09:59:52-0400

a)

"\omega=\omega_0+\alpha t=100-5.56(6)=66.6\frac{rad}{s}"

b)


"\theta=\omega_0t+0.5\alpha t^2\to 500=(100)(6)+0.5\alpha (6)^2"

"\alpha=-5.56\frac{rad}{s^2}"